The utility function for a consumer utility is U=30Q11/2Q21/2 .If the price per unit of Q1 is Ksh 10 and Ksh 5per unit of Q2, determine quantities Q1 and Q2 that the consumer should have to maximize utility if the consumer Ksh 350budgeted
1
Expert's answer
2019-03-01T10:33:08-0500

The condition of optimal consumer choice is:

〖MU〗_1/〖MU〗_2 =P_1/P_2


〖MU〗_1=〖U'〗_1=30×1/2×Q_1^(-1/2)×Q_2^(1/2)

〖MU〗_2=〖U'〗_2=30×1/2×Q_2^(-1/2)×Q_1^(1/2)

〖MU〗_1/〖MU〗_2 =(30×1/2×Q_1^(-1/2)×Q_2^(1/2))/(30×1/2×Q_2^(-1/2)×Q_1^(1/2) )=Q_2/Q_1


Q_2/Q_1 =10/5=2


So, Q2=2Q1

And the budget constraint is 10Q1+5Q2=350

So,

10Q_1+5×2×Q_1=350

20Q_1=350,

Q1=17.5

Q2=35